find r
A=P(1+r)^2
A= 1685.40
P= 1500

Answer by stanbon(75887)

find r
A=P(1+r)^2
A= 1685.40
P= 1500
-----
1685.4 = 1500(1+r)^2
1.1236 = (1+r)^2
1+r = sqrt(1.1236)
1+r = 1.06
r = 0.06
==================
Cheers,
Stan H.